If vectors a, b and c are mutually perpendicular of equal magnitude, then find the angle between vectors a and a+b+c
Please kindly help me solve this.

Answer by Alan3354(69443)

If vectors a, b and c are mutually perpendicular of equal magnitude, then find the angle between vectors a and a+b+c
-------
The 3 vectors form a cube.
The sum of them is parallel to the long diagonal of a cube. If side lengths = s, that's s*sqrt(3).
-------
The cosine of the angle = s*sqrt(2)/(s*sqrt(3)) = sqrt(2/3) = sqrt(6)/3
============================
Lookup "direction cosines"